OpenStudy (anonymous):

From which quadrant line 3/5y-2/5x=0 passes please explain

OpenStudy (yttrium):

As I said, you can multiply it by 5xy \[5xy [\frac{ 3 }{5y } - \frac{ 2 }{ 5x } = 0 ]\] Therefore, your equation will be \[3x - 2y = 0\]

OpenStudy (anonymous):

but from which quadrant will it pass

OpenStudy (yttrium):

Graphing it will.  therefore, from which quadrant will it be?

OpenStudy (anonymous):

@Rohan28 http://www.wolframalpha.com/examples/CoordinateGeometry.html

OpenStudy (anonymous):

\[\frac{ 3\times \sqrt{2}-\sqrt{5} }{ 3+\sqrt{2}-\sqrt{5} }=a+b \times \sqrt{5}\] then find a-b